# What Does It Take to Be a Good AI Research Agent? Studying the Role of Ideation Diversity

### v1

@@ Abstract
Alexis Audran-Reiss at , Jordi Armengol Estapé at
@@ 1 Introduction
The rapid advancement of Large Language Model-based (NEURIPS2020_1457c0d6) agents equipped with tools (schick2023toolformer) has sparked interest in the quest to develop research agents, in areas as challenging as chemistry (Boiko2023) or biology (Swanson2025).
@@
In particular, recent work on autonomous AI research agents (shen2023hugginggptsolvingaitasks; huang2024mlagentbenchevaluatinglanguageagents; toledo2025airesearchagentsmachine; zhao2025automatedllmspeedrunningbenchmark) improves upon automated machine learning engineering tools (autosklearn) by mirroring the cognitive processes of human researchers through a structured research pipeline: idea generation and hypothesis setup, experimental design and implementation, empirical validation, and iterative refinement. Recent advances have achieved notable milestones such as creation of the first fully autonomous AI-generated research paper accepted through peer review (yamada2025aiscientistv2workshoplevelautomated).
Despite the potential of these recent breakthroughs in automating AI science, the field is still in its infancy and little is understood about the factors driving their successes and failures. Error analysis is substantially more complicated than in classic machine learning setups, due to the presence of long multi-step trajectories often guided by heuristic-based search algorithms (toledo2025airesearchagentsmachine) and leveraging tool use, which requires complex evaluation frameworks. Moreover, obtaining large-enough samples to perform meaningful analysis and ablate design choices can be computationally prohibitive.
@@
We perform a first-of-its-kind, large-scale study of AI research agents’ trajectories in MLE-bench (chan2025mlebench), a well-known benchmark of Kaggle machine learning tasks. We study 6 different LLM backbones equipped with 2 different agentic frameworks (or scaffolds) on the 75 machine learning tasks available in MLE-bench across 10 to 20 random seeds, yielding a total of 11,000 trajectories. This corresponds to roughly 1,200,000 individual nodes in the agent scaffold search, for a total of 264,000 GPU hours.
To measure ideation diversity, we propose calculating Shannon entropy (shannon1948mathematical) on the distribution of model architectures that the agent plans to implement in the ideation phase. Figure 1 shows the correlation between ideation diversity and the performance in MLE-bench using our generated trajectory bank, hinting at a relation between the two.
@@ 2 Research Agents and MLE-bench
Broadly, soton252102 define an agent as a computer system that is situated in some environment and that is capable of autonomous action in this environment in order to meet its design objectives. In the particular context of recent work on research agents based on generative AI and LLMs (schick2023toolformer; Boiko2023; shen2023hugginggptsolvingaitasks; huang2024mlagentbenchevaluatinglanguageagents; Swanson2025; toledo2025airesearchagentsmachine; zhao2025automatedllmspeedrunningbenchmark), we specifically refer to agent systems implemented using two main components: 1. a model backbone (typically, an LLM), which processes observations from the environment as prompts and emits text-based actions, and 2. an outer loop making use of the model backbone to interface with the environment . This outer loop orchestrating the LLM actions is usually referred to as agentic frameworks or agentic scaffolds in the literature (wu22). The environments for AI research agents allow the agent to create and run code among other tools (schick2023toolformer).
@@
(chan2025mlebench) (Machine Learning Engineering Benchmark) is a well-known evaluation framework designed to assess autonomous AI agents’ capabilities in solving real-world machine learning problems. The benchmark is constructed from 75 tasks sourced from Kaggle competition datasets, providing a diverse collection of machine learning challenges that span computer vision, natural language processing, time series forecasting, tabular data analysis, and multimodal learning domains.
@@
Given the same dataset, there are multiple ways of defining how to interface with the benchmark, mediated by the agentic scaffold of choice. Recent work (toledo2025airesearchagentsmachine) proposed a tree-based agentic scaffold to tackle MLE-bench. In Section 3.1.3, we describe the scaffolds studied in this work with more detail.
Figure 2 shows an example of a tree-search-based AI research agent attempting the Jigsaw Toxic Comment Classification Challenge MLE-bench task. The AI research agent would execute the following computational workflow: idea generation (e.g., leveraging pre-trained convolutional neural network (lecun98) features with linear classifiers), hypothesis setup (e.g., establishing baselines using CIFAR-100 (krizhevsky2009learning) embeddings with logistic regression), and implementation and experimentation (e.g., tuning learning rate for logistics regression). The next step would be to analyze results (e.g. look at confusion matrix) and finally make a submission to the grader. Based on the leaderboard rank and experiment analysis, the agent would then propose the next set of ideas and hypotheses (e.g. changing classifier from logistic regression to random forest or tuning a CNN end-to-end).
@@ 3.1.3 Agentic Scaffolds
Following recent work (toledo2025airesearchagentsmachine), we formalize AI research agents as search algorithms composed of a search policy, used to navigate the space of candidate solutions to a task, and a set of operators, which modify existing solutions to generate new candidate solutions.
Enabled by this formalization, we study a range of agentic structures, specifically focusing on: (1) AIDE (aide2025), an LLM-driven agent that approaches problem-solving as a tree-search over the domain of Python solutions, utilizing a Greedy policy. (2) $\text{AIRA}_{\text{GREEDY}}$ (toledo2025airesearchagentsmachine), another greedy tree-based search policy, with a different design for operators, memory scope, and prompts, and (3) $\text{AIRA}_{\text{MCTS}}$ (toledo2025airesearchagentsmachine), utilizing Monte Carlo Tree Search (MCTS (mcts1; mcts2; mcts3)) for its search policy, in contrast to its greedy counterparts.
@@ 3.1.4 LLM Backbones
For the data analysis, we use the following LLM backbones for the agents studied: o3 (jaech2024openai), gpt-oss (openai2025gptoss120bgptoss20bmodel) (20B and 120B), Llama Maverick (meta2025llama4), Devstral (rastogi2025devstralfinetuninglanguagemodels) and CWM (cwm2025). Those represent different model sizes and architectures.
We conduct the controlled experiment discussed in Section 3.3 with the full-sized DeepSeek R1 model (deepseekai2025deepseekr1incentivizingreasoningcapability). All of the LLMs above use a 128K-token context window to ensure input coverage without truncation.
@@ 3.2 Measuring Ideation Diversity
All AI research agents in our study begin their exploration by generating at maximum five initial ideas to solve the task at hand (exactly five for greedy searches, and up to five for MCTS), using the Draft operator. To measure ideation diversity, we compare agents by extracting two pieces of information from these five initial ideas. First, we extract the high-level ML approach or architecture used by our agent (for example CNN (lecun98), Transformer (vaswani2017attention), Decision Trees); and second, we also extract the specific model employed by the agent, with variants grouped together (e.g., EfficientNet-B4 is grouped as EfficientNet (tan2019efficientnet)).
@@ 4.2.2 Results
Figure 6: Comparison of MLE-bench lite medal rate of $\text{AIRA}_{\text{Greedy}}$ and $\text{AIRA}_{\text{MCTS}}$ with and without interventions to reduce solution diversity (as indicated by ’- Low diversity’). Error bars represent 95% confidence intervals computed using stratified bootstrapping, using the rliable library (agarwal2021deep).
@@ 4.3.1 Alternative Metrics
Each of the alternative metrics we consider offers a distinct perspective: while some, like medal rate, emphasize marginal improvements, others account for all performance gains. Additionally, certain metrics are based entirely on human score distributions, whereas others operate independently of them. We consider 4 additional metrics. 1. Valid Submission Rate: The percentage of tasks in which the agent is able to make a valid submission. This metric captures the ability of the agent to ideate, implement, and debug until reaching at least one valid submission. 2. Average Normalized Score: For each agent attempt at a task, we compute a normalized score: a score of 0 represents the lowest human score achieved on the task, and 1 the highest. The metric captures how good agents submissions are, independently from human score distributions. 3. Percentile: The metric captures the ability of the agent to outperform humans at machine learning engineering. Compared to medal rate, this metric still relies on human score distributions, and offer a less discrete assessment of the performance. Like average normalized scores, and unlike medal rates, improvements of percentile in poor and strong scores are equally valued. 4. ELO-Based Agent Ranking: We create an ELO system (elo) using all possible heads-to-heads between agents’ scores. ELO rankings are agnostic of the human score distribution on MLE-bench tasks. ELO difference of 100 points corresponds to about a 64% expected win probability for the higher-rated agent.
@@ 4.3.3 Controlled Experiment Results with Additional Metrics
A notable observation is the drop in valid submission rates, which fall from 98% to 92% for $\text{AIRA}_{\text{Greedy}}$ and to 90% for $\text{AIRA}_{\text{MCTS}}$. This indicates that, for certain tasks, Low Diversity agents are sometimes unable to produce even a single valid submission during their search. Our analysis reveals that this decline is primarily driven by two competitions: ’text-normalization-challenge-english-language’ and ’text-normalization-challenge-russian-language’. Upon examining agent trajectories, we find that Low Diversity agents repeatedly attempt to implement the same model, T5 (t5), but consistently fail, resulting in timeouts. In contrast, baseline agents implement a wider range of solutions and are more often able to make correct submissions. Notably, baseline agents also occasionally attempt to implement T5 and encounter similar failures, but their greater ideation diversity allows them to succeed elsewhere. These two competitions also account for a significant portion (estimated at around half) of the observed decline in medal rates.
@@ 5 Discussion
Unsurprisingly, implementation quality is an important bottleneck of AI research agents. We observe a strong correlation between AI research agents’ performance and the ability to implement sufficiently complex solutions. By aggregating performances of AIRA (Greedy and MCTS) for each LLM, Figure 11 shows that, on average, the more time an agent spends on each successfully implemented solution (including ideation, implementation, and model training), the more medals it earns. This suggests that performance increases with the agents’ ability to implement more complex solutions. Furthermore, Figure 11 shows that agents perform better when, out of the 24 hours allotted to complete a task, they spend a higher proportion of time on successfully implemented solutions. However, since LLMs and coding agents are improving rapidly (kwa2025measuringaiabilitycomplete), particularly in verifiable tasks (deepseekai2025deepseekr1incentivizingreasoningcapability), we hypothesize that the relative importance of the ideation and planning phase might increase over time, not to de-risk implementation pitfalls, but to efficiently explore the solution space.
@@ 6 Related Work
Generation diversity in language models has been studied and even explicitly promoted since the statistical machine translation era (macherey-och-2007-empirical; gimpel-etal-2013-systematic; xiao2013bagging), where selecting (devlin-matsoukas-2012-trait) or combining (macherey-och-2007-empirical) a set of diverse yet plausible generations lead to improved translation quality. Similar observations were later identified in neural machine translation and other sequence-to-sequence settings (li2015diversity; vijayakumar2016diverse; ippolito2019comparison). holtzman2020curiouscaseneuraltext developed nucleus sampling with the goal of generating both coherent and diverse text. More recently, Murthy_2025 and DBLP:conf/iclr/KirkMNLHGR24 study the effect of RLHF (ouyang2022traininglanguagemodelsfollow) on LLMs with a focus on the (decreased) generation diversity. chen2024diversitysyntheticdataimpact study effect of diversity of synthetic data in training LLMs. li2025preserving propose a diversity-preserving algorithm for supervised fine-tuning of LLMs.
@@
Trajectory diversity is synonym of increased exploration in reinforcement learning models. hong2018div-driven investigate a diversity-driven exploration strategy for training reinforcement learning models. eysenbach2018diversity propose a diversity objective to learn skills without reward functions. holder2020 improve diversity in population-based reinforcement learning, while conti2018 propose a novelty objective in a population of agents to improve exploration in reinforcement learning. More recently, yao2025diversityawarepolicyoptimizationlarge propose a diversity-aware policy optimization algorithm; unlike the works cited above, it does so in the context of LLMs. zeng2025bstar propose B-star, a reinforcement learning approach for reasoning LLMs that balances exploration and exploitation.
@@
Diversity has also been studied in the context of multi-agent foundation models (tuyls2023). bettini2025impactbehavioraldiversitymultiagent study the impact of behavioral diversity in multi-agent reinforcement learning. li25aamasdiversity address how agents in multi-agent RL often end up learning nearly identical behaviors when they share network parameters, which hurts exploration. The authors propose CTEM, a method that uses trajectory entropy maximization to push agents toward more diverse behaviors without needing complex density models. li2025llmassisted find that generating diverse teammates in multi-agent training can lead to random, semantically meaningless behaviors, reducing training efficiency. Their SemDiv approach leverages large language models to describe coordination strategies in natural language, then converts these into reward functions for training meaningful teammate policies. In LLM-agent world simulations, chu2025exploringcontrollingdiversityllmagent investigate prompt design’s impact on conversational diversity and introduce a prompt-tuning mechanism that controls diversity via a single parameter.
@@
In the era of tool-use LLM-based agents (schick2023toolformer; kaddour2023challengesapplicationslargelanguage), shen2023hugginggptsolvingaitasks; nathani2025mlgymnewframeworkbenchmark evaluate agents at implementing machine learning tasks. MLE-bench (chan2025mlebench), the benchmark used in this work, is a machine learning benchmark consisting of 75 Kaggle tasks. zhao2025automatedllmspeedrunningbenchmark propose a self-contained machine learning task with a focus on language models. AIRA (toledo2025airesearchagentsmachine), studies AI research agents by formalizing AI research as search policies over a space of candidate solutions. This development is parallel to agentic benchmarks and scaffolds for other fields, such as software engineering (jimenez2024swebench; jiang2025aideaidrivenexplorationspace).

### v2

@@ Abstract
Alexis Audran-Reiss at , Jordi Armengol-Estapé at
@@ 1 Introduction
The rapid advancement of Large Language Model-based (Brown et al., 2020) agents equipped with tools (Schick et al., 2023) has sparked interest in the quest to develop research agents, in areas as challenging as chemistry (Boiko et al., 2023) or biology (Swanson et al., 2025).
@@
In particular, recent work on autonomous AI research agents (Shen et al., 2023; Huang et al., 2024; Toledo et al., 2025; Zhao et al., 2025) improves upon automated machine learning engineering tools (Feurer et al., 2022) by mirroring the cognitive processes of human researchers through a structured research pipeline: idea generation and hypothesis setup, experimental design and implementation, empirical validation, and iterative refinement. Recent advances have achieved notable milestones such as creation of the first fully autonomous AI-generated research paper accepted through peer review (Yamada et al., 2025).
Despite the potential of these recent breakthroughs in automating AI science, the field is still in its infancy and little is understood about the factors driving their successes and failures. Error analysis is substantially more complicated than in classic machine learning setups, due to the presence of long multi-step trajectories often guided by heuristic-based search algorithms (Toledo et al., 2025) and leveraging tool use, which requires complex evaluation frameworks. Moreover, obtaining large-enough samples to perform meaningful analysis and ablate design choices can be computationally prohibitive.
@@
We perform a first-of-its-kind, large-scale study of AI research agents’ trajectories in MLE-bench (Chan et al., 2025), a well-known benchmark of Kaggle machine learning tasks. We study 6 different LLM backbones equipped with 2 different agentic frameworks (or scaffolds) on the 75 machine learning tasks available in MLE-bench across 10 to 20 random seeds, yielding a total of 11,000 trajectories. This corresponds to roughly 1,200,000 individual nodes in the agent scaffold search, for a total of 264,000 GPU hours.
To measure ideation diversity, we propose calculating Shannon entropy (Shannon, 1948) on the distribution of model architectures that the agent plans to implement in the ideation phase. Figure 1 shows the correlation between ideation diversity and the performance in MLE-bench using our generated trajectory bank, hinting at a relation between the two.
@@ 2 Research Agents and MLE-bench
Broadly, Wooldridge and Jennings (1995) define an agent as a computer system that is situated in some environment and that is capable of autonomous action in this environment in order to meet its design objectives. In the particular context of recent work on research agents based on generative AI and LLMs (Schick et al., 2023; Boiko et al., 2023; Shen et al., 2023; Huang et al., 2024; Swanson et al., 2025; Toledo et al., 2025; Zhao et al., 2025), we specifically refer to agent systems implemented using two main components: 1. a model backbone (typically, an LLM), which processes observations from the environment as prompts and emits text-based actions, and 2. an outer loop making use of the model backbone to interface with the environment . This outer loop orchestrating the LLM actions is usually referred to as agentic frameworks or agentic scaffolds in the literature (Wu et al., 2022). The environments for AI research agents allow the agent to create and run code among other tools (Schick et al., 2023).
@@
(Chan et al., 2025) (Machine Learning Engineering Benchmark) is a well-known evaluation framework designed to assess autonomous AI agents’ capabilities in solving real-world machine learning problems. The benchmark is constructed from 75 tasks sourced from Kaggle competition datasets, providing a diverse collection of machine learning challenges that span computer vision, natural language processing, time series forecasting, tabular data analysis, and multimodal learning domains.
@@
Given the same dataset, there are multiple ways of defining how to interface with the benchmark, mediated by the agentic scaffold of choice. Recent work (Toledo et al., 2025) proposed a tree-based agentic scaffold to tackle MLE-bench. In Section 3.1.3, we describe the scaffolds studied in this work with more detail.
Figure 2 shows an example of a tree-search-based AI research agent attempting the Jigsaw Toxic Comment Classification Challenge MLE-bench task. The AI research agent would execute the following computational workflow: idea generation (e.g., leveraging pre-trained convolutional neural network (LeCun and Bengio, 1998) features with linear classifiers), hypothesis setup (e.g., establishing baselines using CIFAR-100 (Krizhevsky et al., 2009) embeddings with logistic regression), and implementation and experimentation (e.g., tuning learning rate for logistics regression). The next step would be to analyze results (e.g. look at confusion matrix) and finally make a submission to the grader. Based on the leaderboard rank and experiment analysis, the agent would then propose the next set of ideas and hypotheses (e.g. changing classifier from logistic regression to random forest or tuning a CNN end-to-end).
@@ 3.1.3 Agentic Scaffolds
Following recent work (Toledo et al., 2025), we formalize AI research agents as search algorithms composed of a search policy, used to navigate the space of candidate solutions to a task, and a set of operators, which modify existing solutions to generate new candidate solutions.
Enabled by this formalization, we study a range of agentic structures, specifically focusing on: (1) AIDE (Jiang et al., 2025a), an LLM-driven agent that approaches problem-solving as a tree-search over the domain of Python solutions, utilizing a Greedy policy. (2) $\text{AIRA}_{\text{GREEDY}}$ (Toledo et al., 2025), another greedy tree-based search policy, with a different design for operators, memory scope, and prompts, and (3) $\text{AIRA}_{\text{MCTS}}$ (Toledo et al., 2025), utilizing Monte Carlo Tree Search (MCTS (Coulom, 2006; Kocsis and Szepesvári, 2006; Browne et al., 2012)) for its search policy, in contrast to its greedy counterparts.
@@ 3.1.4 LLM Backbones
For the data analysis, we use the following LLM backbones for the agents studied: o3 (Jaech et al., 2024), gpt-oss (OpenAI, 2025) (20B and 120B), Llama Maverick (Team, 2025b), Devstral (Team, 2025a) and CWM (FAIR CodeGen Team, 2025). Those represent different model sizes and architectures.
We conduct the controlled experiment discussed in Section 3.3 with the full-sized DeepSeek R1 model (DeepSeek-AI et al., 2025). All of the LLMs above use a 128K-token context window to ensure input coverage without truncation.
@@ 3.2 Measuring Ideation Diversity
All AI research agents in our study begin their exploration by generating at maximum five initial ideas to solve the task at hand (exactly five for greedy searches, and up to five for MCTS), using the Draft operator. To measure ideation diversity, we compare agents by extracting two pieces of information from these five initial ideas. First, we extract the high-level ML approach or architecture used by our agent (for example CNN (LeCun and Bengio, 1998), Transformer (Vaswani et al., 2017), Decision Trees); and second, we also extract the specific model employed by the agent, with variants grouped together (e.g., EfficientNet-B4 is grouped as EfficientNet (Tan and Le, 2019)).
@@ 4.2.2 Results
Figure 6: Comparison of MLE-bench lite medal rate of $\text{AIRA}_{\text{Greedy}}$ and $\text{AIRA}_{\text{MCTS}}$ with and without interventions to reduce solution diversity (as indicated by ’- Low diversity’). Error bars represent 95% confidence intervals computed using stratified bootstrapping, using the rliable library (Agarwal et al., 2021).
@@ 4.3.1 Alternative Metrics
Each of the alternative metrics we consider offers a distinct perspective: while some, like medal rate, emphasize marginal improvements, others account for all performance gains. Additionally, certain metrics are based entirely on human score distributions, whereas others operate independently of them. We consider 4 additional metrics. 1. Valid Submission Rate: The percentage of tasks in which the agent is able to make a valid submission. This metric captures the ability of the agent to ideate, implement, and debug until reaching at least one valid submission. 2. Average Normalized Score: For each agent attempt at a task, we compute a normalized score: a score of 0 represents the lowest human score achieved on the task, and 1 the highest. The metric captures how good agents submissions are, independently from human score distributions. 3. Percentile: The metric captures the ability of the agent to outperform humans at machine learning engineering. Compared to medal rate, this metric still relies on human score distributions, and offer a less discrete assessment of the performance. Like average normalized scores, and unlike medal rates, improvements of percentile in poor and strong scores are equally valued. 4. ELO-Based Agent Ranking: We create an ELO system (Bradley and Terry, 1952) using all possible heads-to-heads between agents’ scores. ELO rankings are agnostic of the human score distribution on MLE-bench tasks. ELO difference of 100 points corresponds to about a 64% expected win probability for the higher-rated agent.
@@ 4.3.3 Controlled Experiment Results with Additional Metrics
A notable observation is the drop in valid submission rates, which fall from 98% to 92% for $\text{AIRA}_{\text{Greedy}}$ and to 90% for $\text{AIRA}_{\text{MCTS}}$. This indicates that, for certain tasks, Low Diversity agents are sometimes unable to produce even a single valid submission during their search. Our analysis reveals that this decline is primarily driven by two competitions: ’text-normalization-challenge-english-language’ and ’text-normalization-challenge-russian-language’. Upon examining agent trajectories, we find that Low Diversity agents repeatedly attempt to implement the same model, T5 (Raffel et al., 2020), but consistently fail, resulting in timeouts. In contrast, baseline agents implement a wider range of solutions and are more often able to make correct submissions. Notably, baseline agents also occasionally attempt to implement T5 and encounter similar failures, but their greater ideation diversity allows them to succeed elsewhere. These two competitions also account for a significant portion (estimated at around half) of the observed decline in medal rates.
@@ 5 Discussion
Unsurprisingly, implementation quality is an important bottleneck of AI research agents. We observe a strong correlation between AI research agents’ performance and the ability to implement sufficiently complex solutions. By aggregating performances of AIRA (Greedy and MCTS) for each LLM, Figure 11 shows that, on average, the more time an agent spends on each successfully implemented solution (including ideation, implementation, and model training), the more medals it earns. This suggests that performance increases with the agents’ ability to implement more complex solutions. Furthermore, Figure 11 shows that agents perform better when, out of the 24 hours allotted to complete a task, they spend a higher proportion of time on successfully implemented solutions. However, since LLMs and coding agents are improving rapidly (Kwa et al., 2025), particularly in verifiable tasks (DeepSeek-AI et al., 2025), we hypothesize that the relative importance of the ideation and planning phase might increase over time, not to de-risk implementation pitfalls, but to efficiently explore the solution space.
@@ 6 Related Work
Generation diversity in language models has been studied and even explicitly promoted since the statistical machine translation era (Macherey and Och, 2007; Gimpel et al., 2013; Xiao et al., 2013), where selecting (Devlin and Matsoukas, 2012) or combining (Macherey and Och, 2007) a set of diverse yet plausible generations lead to improved translation quality. Similar observations were later identified in neural machine translation and other sequence-to-sequence settings (Li et al., 2015; Vijayakumar et al., 2016; Ippolito et al., 2019). Holtzman et al. (2020) developed nucleus sampling with the goal of generating both coherent and diverse text. More recently, Murthy et al. (2025) and Kirk et al. (2024) study the effect of RLHF (Ouyang et al., 2022) on LLMs with a focus on the (decreased) generation diversity. Chen et al. (2024) study effect of diversity of synthetic data in training LLMs. Li et al. (2025c) propose a diversity-preserving algorithm for supervised fine-tuning of LLMs.
@@
Trajectory diversity is synonym of increased exploration in reinforcement learning models. Hong et al. (2018) investigate a diversity-driven exploration strategy for training reinforcement learning models. Eysenbach et al. (2019) propose a diversity objective to learn skills without reward functions. Parker-Holder et al. (2020) improve diversity in population-based reinforcement learning, while Conti et al. (2018) propose a novelty objective in a population of agents to improve exploration in reinforcement learning. More recently, Yao et al. (2025) propose a diversity-aware policy optimization algorithm; unlike the works cited above, it does so in the context of LLMs. Zeng et al. (2025) propose B-star, a reinforcement learning approach for reasoning LLMs that balances exploration and exploitation.
@@
Diversity has also been studied in the context of multi-agent foundation models (Tuyls, 2023). Bettini et al. (2025) study the impact of behavioral diversity in multi-agent reinforcement learning. Li and Zhu (2025) address how agents in multi-agent RL often end up learning nearly identical behaviors when they share network parameters, which hurts exploration. The authors propose CTEM, a method that uses trajectory entropy maximization to push agents toward more diverse behaviors without needing complex density models. Li et al. (2025a) find that generating diverse teammates in multi-agent training can lead to random, semantically meaningless behaviors, reducing training efficiency. Their SemDiv approach leverages large language models to describe coordination strategies in natural language, then converts these into reward functions for training meaningful teammate policies. In LLM-agent world simulations, Chu et al. (2025) investigate prompt design’s impact on conversational diversity and introduce a prompt-tuning mechanism that controls diversity via a single parameter.
@@
In the era of tool-use LLM-based agents (Schick et al., 2023; Kaddour et al., 2023), Shen et al. (2023); Nathani et al. (2025) evaluate agents at implementing machine learning tasks. MLE-bench (Chan et al., 2025), the benchmark used in this work, is a machine learning benchmark consisting of 75 Kaggle tasks. Zhao et al. (2025) propose a self-contained machine learning task with a focus on language models. AIRA (Toledo et al., 2025), studies AI research agents by formalizing AI research as search policies over a space of candidate solutions. This development is parallel to agentic benchmarks and scaffolds for other fields, such as software engineering (Jimenez et al., 2024; Jiang et al., 2025b).